A block weighs 100kg and is resting on a plane that is incline to the
horizontal at an angle of 20 degrees
. What is the force that tends to move the block
down the plane?
1
Expert's answer
2021-08-31T08:58:44-0400
"F=mg\sin 20\\F=(100)(9.8)\sin 20=335\ N"